OpenStudy (anonymous):

X^a * X^b * X^c = 1, then a^3+b^3+c^3 = ? a) 9 b) abc c) a+b+c d) 3abc

mathslover (mathslover):

\[\large{X^a*X^b*X^c=1}\] \[\large{X^{a+b+c}=1}\] \[\large{X^{a+b+c}=X^{0}}\] \[\large{a+b+c=0}\]

mathslover (mathslover):

Now can you tell me that what will be a^3+ b^3+c^3 is a+b+c = 0 ?

mathslover (mathslover):

@dhiwagm

OpenStudy (anonymous):

dats what m asking u @mathslover

mathslover (mathslover):

Ok so can you tell me the formula for a^3 + b^3 + c^3 ?

OpenStudy (anonymous):

a^3 + b^3 + c^3 + 3a^2b + 3ab^2 + 3a^c + 3ac^2 + 3b^2c + 3bc^2 + 6abc

mathslover (mathslover):

Ok so it is : \[\large{a^3+b^3+c^3 = (a+b+c)(a^2+b^2+c^2-ab-bc-ca) + 3abc}\] is it this? please check yours' one again

OpenStudy (anonymous):

if a+b+c=0 then a^3 + b^3 + c^3 = 3abc

mathslover (mathslover):

right geedee but let dhiwagm be explained about this...

mathslover (mathslover):

That can only be done if he knows the CORRECT formula for a^3+b^3+c^3

OpenStudy (anonymous):

can you derive d formula plz.. @mathslover

mathslover (mathslover):

We can verify that if you want

Parth (parthkohli):

Actually, no need to waste your time doing what you already know. http://www.wolframalpha.com/input/?i=%28a%2Bb%2Bc%29%28a%5E2%2Bb%5E2%2Bc%5E2%E2%88%92ab%E2%88%92bc%E2%88%92ca%29%2B3abc